Springer, Inc. produces rulers from plastic resin. On March 1, there are 5,000 completed rulers and 5,200 kilograms of resin on hand. Flyer has estimated production and sales of rulers in units for the next 4 months as:

March

April

May

June

Estimated production

26,000

22,000

30,000

32,000

Estimated sales

25,000

21,000

33,000

24,000

Each ruler requires 0.25 kilograms of resin. The cost of resin is $4.40 per kilogram. Flyer wants to have 20% of the next month’s material requirements on hand at the end of each month. Prepare a direct materials purchases budget for May.
Answer #1

Direct material purchase budget for May

Estimated production = 30,000 units

Resin per unit = 0.25 kg

Resin required for production = 7,500 kg

Add: desired ending inventory = 32,000*0.25*20% = 1600 kg

Total required = 9100 kg

Less: beginning inventory = 7500*20% = 1500 kg

Required purchases = 7600 kgs

Cost per kg =$4.40

Budgeted cost =$33,440